OpenStudy (anonymous):

solve the system by the substitution method x-7y=8 2x-14y=11

OpenStudy (anonymous):

x = 8+7y 2(8+7y)-14y = 11 16+14y-14y = 11 ........ i think u maybe made a typo?

OpenStudy (anonymous):

x-7y=8 Rearranging above we get x = 8 + 7y Substituting this value of x in second equation 2(8 + 7y) -14y=11 16 +14y - 14y = 11 This gives 16 = 11 which is wrong Therefore the two given equations do not have a solution.

OpenStudy (anonymous):

In case the answer key gives an answer, then u might hv typed the question wrong

OpenStudy (amistre64):

x-7y=8 2x-14y=11 2x-14y=16 2x-14y=11 which are parallel lines with different constant so they never meet

OpenStudy (anonymous):

thanks amistre!!